OpenStudy (anonymous):

Which equation represents the line passing through the points (3, 2) and (-9, 6)?

OpenStudy (anonymous):

first find the gradient using the formula m = (y2-y1)/(x2-x1) and then find the equation using the point-gradient formula y - y1 = m(x - x1).